A study was conducted to determine the effectiveness of a weight loss program. The table below shows the before and after weights of 10 subjects in the program. Is this program effective for reducing weights? (alpha = 0.05)

Participant Before After

1 185 169

2 192 187

3 206 193

4 177 176

5 225 194

6 168 171

7 256 228

8 239 217

9 199 204

10 218 195

QUESTION:

a) What are the critical values?

b) Calculate the t value
Answer #1

a) for a = 0.05 and d.f = n -1 = 10 -1 = 9

tCritical = t​​​​​​0.05, 9

tCritical = -2.228

b) test statistic

t = dbar / sd/√n

t = -13.1/4.12

t = -3.18